# OCRBench v2: An Improved Benchmark for Evaluating Large Multimodal Models on Visual Text Localization and Reasoning

### v1

@@ 3.2 Annotation Curation
Qwen-VL [5] 34.6 7.5 0 18.2 20.0 8.1 57.2 41.1 23.3
Qwen-VL-chat [5] 34.5 4.1 0 25.9 14.0 13.8 55.7 39.5 23.4
Qwen2-VL-8B [105] 72.1 47.9 17.5 82.5 25.5 25.4 78.4 61.5 51.4
InternVL2-8B [13] 49.9 23.1 0.5 65.2 24.8 26.7 73.5 52.9 39.6
InternVL2-26B [13] 63.4 26.1 0 76.8 37.8 32.3 79.4 58.9 46.8
InternVL2.5-8B [12] 59.0 25.0 1.4 77.5 35.1 29.4 75.3 57.2 45.0
InternVL2.5-26B [12] 65.6 26.1 1.6 86.9 36.2 37.4 78.3 62.9 49.4
@@
EMU2-chat [95] 42.1 0.2 0 12.5 8.1 11.2 42.7 33.4 18.8
mPLUG-Owl3 [120] 41.6 14.0 0.6 24.4 10.9 11.1 52.2 46.0 25.1
CogVLM-chat [106] 50.9 0 0 0.2 8.4 15.0 58.1 41.7 21.8
Deepseek-VL-7B [62] 37.1 15.4 0 23.5 14.6 20.8 53.3 52.9 27.2
@@
MiniCPM-V-2.6 [117] 66.8 6.0 0.8 62.0 28.8 32.4 73.7 52.1 40.3
TextHarmony [133] 25.8 2.5 0 1.8 8.5 10.4 46.1 33.1 16.0
VILA1.5-8B [60] 35.3 15.5 0 21.1 12.7 17.3 46.3 40.3 23.6
LLaVAR [131] 37.3 0 0 1.0 9.9 12.3 34.6 27.0 15.3
DocOwl2 [32] 24.0 9.7 0 13.4 13.5 8.8 53.7 32.0 19.4
UReader [119] 22.4 0.1 0 0 9.2 7.9 41.0 29.1 13.7
Yi-VL-6B [122] 28.9 2.9 0 9.7 12.9 15.8 36.1 32.0 17.3
Janus-1.3B [110] 46.1 0 0 0.2 14.5 13.5 36.0 39.1 18.7
Cambrian-1-8B [100] 45.3 21.5 0 53.6 19.2 19.5 63.5 55.5 34.7
LLaVA-OV-7B [47] 46.0 20.8 0.1 58.3 25.3 23.3 64.4 53.0 36.4
Eagle-X5-7B [92] 34.7 17.8 0 21.7 20.6 21.5 61.0 42.6 27.5
Idefics3-8B [44] 23.8 13.2 0 63.2 23.8 23.0 65.8 44.9 32.2
Ovis1.6-3B [66] 59.2 14.3 0 65.0 32.1 29.0 69.8 56.8 40.8
Pixtral-12B [2] 48.9 21.6 0 66.3 35.5 29.8 66.9 53.7 40.3
@@
Table 2: Evaluation of existing LMMs on English subsets of OCRBench v2. “Recognition” refers to text recognition, “Referring” to text referring, “Spotting” to text spotting, “Extraction” to relation extraction, “Parsing” to element parsing, “Calculation” to mathematical calculation, “Understanding” to visual text understanding, and “Reasoning” to knowledge reasoning. Higher values indicate better performance. Best performance is in boldface, and the second best is underlined. The notations apply to all subsequent figures.
Method Recognition Extraction Parsing Understanding Reasoning Average
Open-source LMMs
Qwen-VL [5] 7.2 5.3 10.7 11.5 11.2 9.2
Qwen-VL-chat [5] 9.5 8.2 9.3 11.0 21.1 11.8
Qwen2-VL-8B [105] 51.3 51.4 21.6 52.5 37.5 42.9
InternVL2-8B [13] 20.6 45.2 23.2 54.4 38.1 36.3
InternVL2-26B [13] 21.9 46.0 34.8 50.9 34.8 37.7
InternVL2.5-8B [12] 52.8 52.8 28.6 56.4 40.5 46.2
InternVL2.5-26B [12] 32.4 56.1 32.6 56.3 43.6 44.2
TextMonkey [59] 23.5 14.8 8.4 19.9 12.2 15.8
LLaVA-Next-8B [52] 5.7 2.9 12.2 7.5 17.2 9.1
Monkey [48] 4.6 11.2 8.4 21.5 20.0 13.1
XComposer2-4KHD [21] 16.7 18.8 12.1 27.5 2.3 15.5
Molmo-7B [17] 7.1 15.0 9.2 9.0 23.7 12.8
EMU2-chat [95] 2.3 0.5 8.5 1.0 7.3 3.9
mPLUG-Owl3 [120] 6.6 17.9 9.7 6.0 26.1 13.3
CogVLM-chat [106] 5.5 10.0 9.8 1.5 2.5 5.9
Deepseek-VL-7B [62] 8.0 13.3 15.7 5.5 18.5 12.2
GLM-4V-9B [28] 24.4 60.6 20.4 52.8 25.2 36.6
MiniCPM-V-2.6 [117] 51.0 29.9 21.2 34.0 33.6 33.9
TextHarmony [133] 1.8 4.5 8.2 1.5 11.9 5.6
VILA1.5-8B [60] 5.4 8.8 8.5 3.0 15.5 8.2
LLaVAR [131] 2.3 1.7 8.9 0 2.5 3.1
DocOwl2 [32] 4.2 10.3 8.6 4.0 9.6 7.3
UReader [119] 6.8 2.7 8.4 2.5 7.2 5.5
Yi-VL-6B [122] 4.8 4.4 8.5 4.0 25.0 9.4
Janus-1.3B [110] 7.6 8.7 11.4 4.5 10.7 8.6
Cambrian-1-8B [100] 5.3 14.9 12.6 8.5 8.1 9.9
LLaVA-OV-7B [47] 14.8 15.7 13.7 16.0 28.7 17.8
Eagle-X5-7B [92] 7.5 12.0 11.6 5.0 19.2 11.1
Idefics3-8B [44] 7.0 15.5 15.9 9.0 18.1 13.1
Ovis1.6-3B [66] 11.5 23.7 22.8 28.8 18.9 21.1
Pixtral-12B [2] 13.4 10.9 21.0 7.0 20.7 14.6
Closed-source LMMs
GLM-4V-Plus [28] 34.5 60.6 23.9 49.8 28.2 39.4
GPT-4V [78] 49.9 52.2 34.6 40.8 22.9 40.1
GPT-4o [1] 21.6 53.0 29.8 38.5 18.2 32.2
GPT-4o-mini [77] 13.1 38.9 27.2 28.8 16.9 25.0
Gemini-Pro [99] 52.5 47.3 30.9 51.5 33.4 43.1
Claude3.5-sonnet [3] 21.0 56.2 35.2 55.0 30.5 39.6
Step-1V [94] 56.7 41.1 37.6 38.3 39.2 42.6
Table 3: Evaluation of existing LMMs on Chinese subsets of OCRBench v2.
3.3 Statistics of OCRBench v2
Here we present the statistics of OCRBench v2, including the proportion of eight task categories, the OCR-related statistics, and the measurement of prompt quality. More details on statistics can be found in Appendix D.
Capabilities Distribution. The proportion of eight capabilities in OCRBench v2 is shown in Figure 4. The proportions across different categories are relatively balanced.
OCR-Related Statistics. We further count the distribution of line-level OCR results of $7,400$ English and $2,600$ Chinese images in Figure 5(a) and Figure 5(b), respectively. The average number of line-level OCR results per category is shown in Figure 5(c). These statistics demonstrate that the text information is sufficiently rich in OCRBench v2.
Prompt Quality. We also compare the metrics of Average Entropy, Type-Token Ratio, and Average Variability Index of the questions between OCRBench v2 and OCRBench in Figure 6. Compared with OCRBench, OCRBench v2 shows greater Average Entropy, indicating higher unpredictability and diversity in its questions, which challenges the LMMs to tackle more diverse scenarios. It also exhibits a higher Type-Token Ratio, reflecting greater lexical variety and less redundancy in its question formulation. Additionally, the Average Variability Index of OCRBench v2 is significantly higher, suggesting a wider range of question types and structures, offering a more thorough performance evaluation of LMMs.
@@ 3.4 Evaluation Criteria
Since OCRBench v2 contains $23$ distinct tasks, a single evaluation metric is inadequate. Therefore, we adopt six types of evaluation metrics tailored to specific task categories. In the following, we present an overview of the evaluation metrics and their applicability to specific tasks. More details can be found in Appendix E.
@@
Long Reading Type. To assess performance on tasks involving long text reading, BLEU [80], METEOR [6], F1 score, and edit distance are used to assess the similarity between predicted text and ground truth.
@@
Basic VQA Type. To evaluate LMMs on general VQA cases, we employ task-specific metrics based on answer length and format. For questions where the original data provide options, we use exact string matching to compute accuracy. In other cases, we follow the approach of OCRBench to check whether the ground truth is contained in the prediction for short answers (fewer than $5$ words) and employ ANLS to measure prediction quality for longer answers ($5$ words or more).
4 Results and Discussions
Method Resolition Recognition Referring Spotting Extraction Parsing Calculation Understanding Reasoning Average
InternVL2-8B [13] 448 47.3 19.1 0.1 52.8 27.3 25.4 61.1 49.1 35.3
896 48.7 23.0 0.5 66.2 26.2 25.9 73.2 51.9 39.4
dynamic 49.9 23.1 0.5 65.2 24.8 26.7 73.5 52.9 39.6
Table 4: Evaluation of InternVL2-8B with different resolution setting on the English subsets of OCRBench v2.
@@ 4 Results and Discussions
Qwen2VL-8B [105] 72.1 47.9 17.5 82.5 25.5 25.4 78.4 61.5 51.4
Qwen2VL-8B+OCR 69.8 50.4 20.1 79.1 29.4 28.0 77.7 60.0 51.8
Qwen2.5-8B+OCR 28.6 13.8 0 45.9 24.2 31.3 61.1 40.5 30.7
Table 5: Evaluation of Qwen2VL-8B and Qwen2.5-8B with pre-provided OCR information on English subsets of OCRBench v2.
Here we first benchmark state-of-the-art LMMs on OCRBench v2, presenting the quantitative analysis of their OCR capabilities. We then investigate several critical factors influencing OCR performance, including the resolution setting of visual encoder, pre-provided OCR information, and the direct integration between OCR information and language models. Further, we summarize key findings of current limitations and challenges for LMMs. All results are presented as percentages.
Figure 7: Results of eight representative LMMs across the eight ability dimensions defined in OCRBench v2.
4.1 Baselines
We evaluate a total of $31$ open-source LMMs including InternVL series [13, 12], Qwen-VL series [5, 105], GLM-4v-9B [28], Monkey [48], TextMonkey [59], LLaVA-Next-8B [52], XComposer2-4KHD [21], EMU2-chat [95], mPLUG-Owl3 [120], CogVLM-chat [106], Deepseek-VL-7B [62], Molmo-7B [17], MiniCPM-V-2.6 [117], TextHarmony [133], VILA1.5-8B [60], LLaVAR [131], DocOwl2 [32], UReader [119], Yi-VL-6B [122], Janus-1.3B [110], Cambrian-1-8B [100], LLaVA-OV-7B [47], Eagle-X5-7B [92], Idefics3-8B [44], Ovis1.6-3B [66], and Pixtral-12B [2] with the publicly available checkpoints. Further, some cutting-edge commercial models are also tested, including GPT4V [78], GPT4o [79], GPT4o-mini [77], Gemini-Pro [99], Claude3.5-sonnet [3], GLM-4V-Plus [28], and Step-1V [94].
4.2 Main Results
The main evaluation results are shown in Table 2, Table 3, and Figure 7. We observe that, while current LMMs perform relatively well on some basic capabilities such as text recognition and visual text understanding, most LMMs achieve low scores in other capabilities, such as text spotting and element parsing, mostly below $50$. In particular, some LMMs show significant limitations in text spotting capabilities, failing to precisely locate and recognize the texts. Additionally, LMMs demonstrate inadequate abilities in element parsing and mathematical calculation, which are crucial for complicated tasks like document analysis and mathematical reasoning. Besides, after comparing the performance of LMMs on visual text understanding and knowledge reasoning capabilities, we find that they perform poorly in knowledge reasoning, despite being provided with multiple-choice options. This suggests the deficiency of LMMs in logical reasoning.
4.3 Potential Factors Affecting OCR Capabilities
To provide additional insights, we also conduct experiments to explore the potential factors that may affect the performance of LMMs on OCRBench v2.
High-Res Visual Encoders. Since text often appears small in images, the resolution setting of the visual encoder could be a key factor affecting the text perception ability of LMMs [48]. Here we change the input resolution of the LMMs and observe the performance changes. In particular, InternVL2-8B is chosen, and the resolution setting includes $448$, $896$, and dynamic. Table 4 lists the results. Indeed, when the input resolution increases from $448$ to $896$, the performance increases by $4.1\%$.
Pre-provided OCR Information. To study the impact of OCR information, we use PaddleOCR [^2] https://github.com/PaddlePaddle/PaddleOCR to pre-extract OCR results and incorporate them with prompts. Table 5 shows the results. We observe that adding OCR information does not help much. This suggests that OCRBench v2 evaluates LMMs capabilities across multiple dimensions, rather than solely focusing on text recognition abilities.
Connection Between OCR and LLMs. We further explore a direct pipeline by first extracting OCR information and then by feeding it directly into Qwen2.5. Unlike LMMs, this pipeline separates OCR and language modeling into distinct stages. The results shown in Table 5 suggests that Qwen2VL-8B outperforms Qwen2.5 with OCR information, demonstrating LMMs’ remarkable ability to incorporate both textual and visual features efficiently.
@@ 4.4 Main Findings
Here we further highlight some interesting findings from the experimental results:
Finding 1. Limited recognition on less frequently encountered texts. Despite LMMs achieving high performance on text recognition, they can perform poorly when facing with less frequently encountered texts, such as dot matrix texts and mathematical formulas. This performance gap highlights the continuing challenges LMMs face in real-world text recognition.
Finding 2. Limited fine-grained spatial perception. Given that both closed- and open-source LMMs perform poorly in tasks such as text referring and text spotting in OCRBench v2, the current LMMs still lack the capability for accurate text localization.
Finding 3. Insufficient layout perception. While LMMs achieve good performance on basic text recognition, they struggle with complex layouts such as overlapping or rotated texts. For example, GPT-4o fails to detect all characters in overlapping handwritten text and misrecognizes numbers in 90° rotated images, revealing LMMs’ limitations in handling texts with complex layouts.
Finding 4. Weak analytics for complex elements. When facing complex visual elements like charts and formulas, LMMs struggle to parse them into structured data formats, hindering their effectiveness in downstream applications such as document digitalization.
Finding 5. Limited logical reasoning. While LMMs demonstrate basic multimodal reasoning abilities, they struggle with complex scenarios involving mathematical problems or complicated textual reasoning, highlighting the room for improvement in their reasoning capabilities.

### v2

@@ 4.3 Statistics of OCRBench v2
LLaVA-Next-8B [49] 41.3 18.8 0 49.5 21.2 17.3 55.2 48.9 31.5
LLaVA-OV-7B [50] 46.0 20.8 0.1 58.3 25.3 23.3 64.4 53.0 36.4
Monkey [51] 35.2 0 0 16.6 16.3 14.4 59.8 42.3 23.1
@@
Cambrian-1-8B [53] 45.3 21.5 0 53.6 19.2 19.5 63.5 55.5 34.7
Pixtral-12B [54] 48.9 21.6 0 66.3 35.5 29.8 66.9 53.7 40.3
Qwen2.5-VL-7B [21] 68.8 25.7 1.2 80.2 30.4 38.2 73.2 56.2 46.7
InternVL3-14B [23] 67.3 36.9 11.2 89.0 38.4 38.4 79.2 60.5 52.6
Deepseek-VL2-Small [55] 62.7 28.0 0.1 77.5 32.7 14.3 77.1 53.9 43.3
MiniCPM-o-2.6 [56] 66.9 29.5 0.5 70.8 33.4 31.9 69.9 57.9 45.1
@@
Ovis2-8B [58] 73.2 24.6 0.7 62.4 44.8 40.6 72.7 62.6 47.7
@@
Table 2: Evaluation of existing LMMs on English tasks of OCRBench v2’s public data. “Recognition”, “Referring”, “Spotting”, “Extraction”, “Parsing”, “Calculation”, “Understanding”, and “Reasoning” refer to text recognition, text referring, text spotting, relation extraction, element parsing, mathematical calculation, visual text understanding, and knowledge reasoning, respectively. Higher values indicate better performance. Best performance is in boldface, and the second best is underlined. The notations apply to all subsequent figures.
@@ 4.4 Evaluation Criteria
We adopt six types of evaluation metrics tailored to specific task categories. In the following, we present an overview of the evaluation metrics and their applicability to specific tasks.
@@
Long Reading Type. To assess performance on long text reading tasks, BLEU [64], METEOR [65], F1 score, and edit distance are used to assess the similarity between predicted text and ground truth.
@@
Basic VQA Type. For questions where the original data provides options, we use exact string matching to compute accuracy. In other cases, we follow the approach of OCRBench to check whether the ground truth is contained in the prediction for short answers (fewer than $5$ words) and employ ANLS to measure prediction quality for longer answers ($5$ words or more).
5 Results and Findings
Method LLM Size Recognition Extraction Parsing Understanding Reasoning Average
Open-source LMMs
LLaVA-Next-8B [49] 8B 5.7 2.9 12.2 7.5 17.2 9.1
LLaVA-OV-7B [50] 8B 14.8 15.7 13.7 16.0 28.7 17.8
Monkey [51] 8B 4.6 11.2 8.4 21.5 20.0 13.1
TextMonkey [7] 8B 23.5 14.8 8.4 19.9 12.2 15.8
Molmo-7B [52] 8B 7.1 15.0 9.2 9.0 23.7 12.8
Cambrian-1-8B [53] 8B 5.3 14.9 12.6 8.5 8.1 9.9
Pixtral-12B [54] 12B 13.4 10.9 21.0 7.0 20.7 14.6
Qwen2.5-VL-7B [21] 8B 75.3 61.4 41.8 59.3 40.4 55.6
InternVL3-14B [23] 14B 66.2 64.8 33.5 63.4 50.6 55.7
Deepseek-VL2-Small [55] 16B 60.9 50.6 28.3 53.0 20.5 42.7
MiniCPM-o-2.6 [56] 7B 53.0 49.4 27.1 43.5 32.7 41.1
GLM-4V-9B [57] 9B 24.4 60.6 20.4 52.8 25.2 36.6
Ovis2-8B [58] 7B 72.2 50.8 37.7 47.9 37.4 49.2
Closed-source LMMs
GPT-4o [1] - 21.6 53.0 29.8 38.5 18.2 32.2
GPT-4o-mini [59] - 13.1 38.9 27.2 28.8 16.9 25.0
Gemini-Pro [60] - 52.5 47.3 30.9 51.5 33.4 43.1
Claude3.5-sonnet [61] - 21.0 56.2 35.2 55.0 30.5 39.6
Step-1V [62] - 56.7 41.1 37.6 38.3 39.2 42.6
Table 3: Evaluation of existing LMMs on Chinese tasks of OCRBench v2’ public data. “LLM Size” indicates the number of parameters of the language model employed in each method.
Here we first benchmark state-of-the-art LMMs on OCRBench v2, presenting the quantitative analysis, then summarize key findings of current limitations for LMMs. All results are presented as percentages.
5.1 Baselines
The tested LMMs in the section includes LLaVA-Next-8B [49], LLaVA-OV-7B [50], Monkey [51], TextMonkey [7], Molmo-7B [52], Cambrian-1-8B [53], Pixtral-12B [54], Qwen2.5-VL-7B [21], InternVL3-14B [23], Deepseek-VL2-Tiny [55], MiniCPM-o-2.6 [56], GLM-4v-9B [57], Ovis2-8B [58], GPT4o [24], GPT4o-mini [59], Gemini-1.5-Pro [60], Claude3.5-sonnet [61], and Step-1V [62]. Due to space limitations, more LMM evaluation results can be found in the Sec. A.8.
5.2 Main Results
Evaluation results on public data are shown in Tab. 2 and Tab. 3. While LMMs perform well on some basic capabilities such as text recognition and visual text understanding, most LMMs achieve low scores in other capabilities, such as text spotting and element parsing, mostly below $50$. In particular, some LMMs show significant limitations in text spotting capabilities, failing to precisely locate and recognize the texts. Additionally, LMMs demonstrate inadequate abilities in element parsing and mathematical calculation, which are crucial for complicated tasks like document analysis and mathematical reasoning. Besides, after comparing the performance of LMMs on visual text understanding and knowledge reasoning capabilities, we find that they perform poorly in knowledge reasoning. This suggests the deficiency of LMMs in logical reasoning.
Evaluation results on private data are shown in Tab. 4 and Tab. 5. We observe similar evaluation trends to those in the public test set experiments. Overall, LMMs exhibit unsatisfactory performance in text referring, text spotting, element parsing, mathematical calculation, and knowledge reasoning capabilities. In addition, closed-source LMMs outperform their open-source counterparts, demonstrating stronger generalization capabilities. The consistent results across both public and private test sets confirm the soundness of OCRBench v2’s task design, data collection process, and evaluation metrics, and demonstrate its effectiveness in revealing the capability limitations of current LMMs.
@@ 5.2 Main Results
Open-source LMMs
LLaVA-Next-8B [49] 41.4 17.0 0 49.0 12.9 16.1 60.9 30.5 28.5
LLaVA-OV-7B [50] 45.4 18.5 0 60.0 15.5 32.0 59.0 39.3 33.7
Monkey [51] 31.5 0.1 0 34.4 26.3 17.7 61.4 22.4 24.2
TextMonkey [7] 39.8 1.6 0 27.6 24.8 10.2 62.3 21.2 23.4
Molmo-7B [52] 40.8 19.5 0 51.7 10.0 33.9 67.0 48.0 33.9
Cambrian-1-8B [53] 44.0 19.0 0 52.3 19.0 20.7 64.0 39.3 32.3
Pixtral-12B [54] 45.1 21.8 0 71.6 21.7 30.4 77.3 39.5 38.4
Qwen2.5-VL-7B [66] 51.5 24.5 3.1 64.8 13.1 53.3 78.6 45.5 41.8
InternVL3-14B [23] 55.8 24.5 2.1 89.3 21.0 59.5 72.0 50.0 46.8
Deepseek-VL2-Small [55] 56.6 23.7 0 86.4 18.9 30.6 72.2 39.5 41.0
MiniCPM-o-2.6 [56] 54.1 24.7 0.3 74.4 17.6 39.2 75.7 47.0 41.6
GLM-4v-9B [57] 52.7 20.6 0 79.4 15.9 21.5 74.7 32.0 37.1
Ovis2-8B [58] 54.2 20.9 0 83.6 24.2 54.7 74.1 57.3 46.1
Closed-source LMMs
GPT-4o [1] 58.6 23.4 0 87.4 23.1 51.6 74.4 62.3 47.6
GPT-4o-mini [59] 55.3 21.8 0 85.4 20.6 45.2 75.5 49.0 44.1
Gemini1.5-Pro [60] 59.1 41.2 6.6 89.5 22.4 54.7 78.8 60.3 51.6
Claude3.5-sonnet [61] 52.9 24.9 2.5 86.9 23.8 61.4 74.4 53.0 47.5
Step-1V [62] 56.7 27.4 2.6 86.3 33.3 42.6 76.6 48.7 46.8
Table 4: Evaluation of existing LMMs on English tasks of OCRBench v2’s private data.
Method LLM Size Recognition Extraction Parsing Understanding Reasoning Average
Open-source LMMs
LLaVA-Next-8B [49] 8B 2.8 0.9 14.9 20.0 7.4 9.2
LLaVA-OV-7B [50] 8B 5.4 13.6 20.3 34.0 13.6 17.4
Monkey [51] 8B 1.5 28.4 29.1 40.0 8.3 21.5
TextMonkey [7] 8B 10.5 15.2 30.2 44.0 7.6 21.5
Molmo-7B [52] 8B 3.4 29.8 6.6 24.0 11.1 15.0
Cambrian-1-8B [53] 8B 2.4 19.8 26.7 36.0 7.6 18.5
Pixtral-12B [54] 12B 6.2 22.3 11.4 26.0 14.0 16.0
Qwen2.5-VL-7B [66] 8B 24.4 78.9 33.1 82.0 29.0 49.5
InternVL3-14B [23] 14B 62.1 59.5 33.2 80.0 29.2 52.8
DeepSeek-VL2-Small [55] 16B 51.6 56.3 27.8 79.6 25.3 48.1
MiniCPM-o-2.6 [56] 7B 54.0 62.4 24.1 68.0 29.8 47.7
GLM-4v-9B [57] 9B 60.6 65.2 32.4 82.0 18.2 51.7
Ovis2-8B [58] 7B 61.0 67.7 43.6 82.0 25.6 56.0
Closed-source LMMs
GPT-4o [1] - 41.7 52.1 29.0 76.0 29.4 45.7
GPT-4o-mini [59] - 20.0 53.6 27.9 66.0 19.6 37.4
Gemini1.5-Pro [60] - 71.4 63.8 30.5 82.0 29.9 55.5
Claude3.5-sonnet [61] - 34.2 62.5 35.2 78.0 32.2 48.4
Step-1V [62] - 65.2 64.9 33.1 78.0 25.5 53.4
Table 5: Evaluation of existing LMMs on Chinese tasks of OCRBench v2’s private data.
@@ 5.3 Main Findings
We provide in-depth analyses for LMMs’ common limitations, including rare text recognition, fine-grained spatial perception, layout perception, complex element analysis, and logical reasoning.
Finding 1. LMMs still face challenges with less frequently encountered texts, such as dot matrix texts and mathematical formulas. This performance gap highlights the continuing challenges LMMs face in real-world text recognition. For instance, occluded text, CAPTCHA, and dot-matrix text are considered low-frequency text, whereas other types belong to high-frequency text. Notably, recognition accuracy varies significantly across these categories. For example, InternVL3-14B achieves 79.1% accuracy on high-frequency texts but drops to 46.7% on low-frequency ones.
Finding 2. Current LMMs still exhibit limited performance in tasks requiring precise spatial understanding, such as text referring and text spotting. For instance, while InternVL3-14B achieves a response accuracy of 78.3% on the VQA with position task, its IoU score for answer region localization is only 12.9%. This suggests that although LMMs can roughly identify where the answer is located, they struggle to output the exact region.
Finding 3. While LMMs achieve good performance on basic text recognition, they struggle with complex layouts such as overlapping or rotated texts. For example, GPT-4o fails to detect the characters in overlapping handwritten text and misrecognizes numbers in 90° rotated images, revealing LMMs’ limitations in handling texts with complex layouts. Rotating images in the DocVQA dataset led to a significant performance drop of $55.7\%$ for InternVL3-14B (from $90.9\%$ to $35.2\%$).
Finding 4. LMMs still struggle to parse text into structured formats in downstream applications such as document digitalization. For instance, InternVL3-14B achieves 94.4% accuracy in unpaired entities matching, but its performance drops to 84.9% in key information extraction, where the model is required to identify the corresponding value given an entity. The performance further degrades in element parsing tasks that demand structured outputs.
Finding 5. Despite recent advances, LMMs still face challenges in complex mathematical and textual reasoning tasks. To assess their capabilities, we evaluated InternVL3-14B on the private test set covering reasoning VQA, ScienceQA, and APP agent tasks. Questions were categorized into five types: common sense reasoning, visual-text understanding, pattern recognition, calculation, and expert knowledge. Human ratings showed the model achieved accuracies of $72.9\%$, $83.0\%$, $69.2\%$, $56.5\%$, and $71.8\%$, respectively, indicating notable variation.